HUBBLE SPACE TELESCOPE DAILY REPORT #2131 PERIOD COVERED: 0000Z (UTC) 05/26/98 - 0000Z (UTC) 05/27/98 Daily Status Report as of 147/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of NIC/2 7330 (The Fueling of Active Nucleii: A NICMOS Snapshot Survey) The Near Infrared Camera (Camera Two) was used to probe the dark matter distribution of NGC-7130 and NGC-4968 far from their centers by observing intervening absorption from the MgII doublet (lambda 2800) in the spectrum of one background QSO located in a geometrically optimal position. HST observations are required because this feature is in a wavelength range not accessible from the ground, and the QSO magnitudes and expected strength of absorption make IUE observations unfeasible. The proposal completed as expected with no anomalies. 1.2 Completed Two Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed NIC/1/2/3 7372 (Fundamental Plane, Morphology-Density Relation, and Lensing in the z=0.83 Cluster) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of galactic cluster Gal-Clus-1. These data will be used to augment the HST mosaics of the clusters with extensive ground-based spectroscopy and imaging at the Keck, MMT, NTT, and WHT telescopes. The observations were completed as planned, and no problems were reported. 1.5 Completed Two Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.6 Completed Two Sets of NIC/1/2 7956 (NICMOS Camera 1 Narrow Band Filter Pointed Super Flats) The Near Infrared Camera (Cameras One and Two) was used to obtain sequences of NICMOS Camera One narrow band filter flat fields. The observations will be used to build high S/N NICMOS flat fields for the different filters. The observations were completed as planned, and no problems were noted. 1.7 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 9 Successful : 9 2.3 Operations Notes: Using ROP SR-1, the EDAC error counts were cleared two times. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH